# Learning Only with Images: Visual Reinforcement Learning with Reasoning, Rendering, and Visual Feedback

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) exhibit impressive performance across various visual tasks. Subsequent investigations into enhancing their visual reasoning abilities have significantly expanded their performance envelope. However, a critical bottleneck in the advancement of MLLMs toward deep visual reasoning is their heavy reliance on curated image-text supervision. To solve this problem, we introduce a novel framework termed “Reasoning-Rendering-Visual-Feedback” (RRVF), which enables MLLMs to learn complex visual reasoning from only raw images. This framework builds on the “Asymmetry of Verification” principle to train MLLMs, i.e., verifying the rendered output against a source image is easier than generating it. We demonstrate that this relative ease provides an ideal reward signal for optimization via Reinforcement Learning (RL) training, reducing reliance on the image-text supervision. Guided by the above principle, RRVF implements a closed-loop iterative process encompassing reasoning, rendering, and visual feedback components, enabling the model to perform self-correction through multi-turn interactions, while this pipeline can be optimized end-to-end by the GRPO algorithm. Extensive evaluations are conducted on image-to-code generation across two diverse domains: data charts and web interfaces. The RRVF-trained model not only outperforms existing open-source MLLMs and supervised fine-tuning baselines but also exhibits superior generalization to unseen datasets. Critically, the model’s performance surpasses that of the more advanced MLLM used to provide the feedback signal during training. This work establishes a self-improvement paradigm that offers a viable path to robust, generalizable models without reliance on explicit supervision. Code will be available at https://github.com/L-O-I/RRVF.
@@ Introduction
Figure 1: Comparison of training paradigms. Conventional MLLMs require paired image-text data for supervised training. The proposed RRVF framework is trained solely on raw images, utilizing a closed-loop process of reasoning, rendering, and visual feedback that is optimized with the GRPO.
Figure 2: An overview of the Reasoning-Rendering-Visual-Feedback (RRVF) framework. The framework is designed to improve visual reasoning using only image inputs. (a) Given an input image, (b) the multimodal language model engages in an iterative reasoning process to generate rendering code. The generated code is executed by external tools to produce a rendered image. (c) The rendered image is compared to the original, and the discrepancies are converted into structured visual feedback. This feedback guides the model in the subsequent reasoning turn. (d) Upon completion of the iterative process, a visual judge quantitatively assesses the similarity between the final rendered output and the original image. (e) This similarity score, along with rewards for format validity and tool usage, is integrated by a reward manager to compute a final reward signal. This signal is used to optimize the model policy via reinforcement learning with the GRPO algorithm.
@@
Specifically, for MLLMs, effective reasoning necessitates capabilities that extend beyond elementary visual perception to achieve sophisticated visual comprehension and inference. This paradigm requires models to not only perform object recognition but also to analyze underlying visual semantics, spatial-geometric relationships, and complex inter-object dependencies (Su et al. 2025c; Luo et al. 2024). While accurate visual perception serves as a fundamental prerequisite for such reasoning capabilities, it inherently depends on robust reasoning mechanisms, creating a bidirectional dependency.
@@
Moreover, the GRPO optimization is guided by a hybrid reward function. (1) The primary component is a visual similarity reward that quantifies the fidelity between the rendered image and the original image by integrating signals from established vision-language models. (2) This is supplemented by a format correctness reward to ensure the generation of syntactically correct and executable code. And (3) an adaptive tool-use efficiency reward is introduced to balance exploration and convergence. This hybrid reward function effectively enables the model to learn underlying generative logic directly from pixel-level visual representations.
@@
  • The exploration of the ”Asymmetry of Verification” principle for MLLM training in complex visual reasoning tasks. We establish that verifying the visual correctness of a rendered output against a source image is a significantly more tractable problem than generating the initial code. This asymmetry provides an effective learning signal that circumvents the need for textual supervision.
  • A novel framework, termed Reasoning-Rendering-Visual-Feedback (RRVF), is proposed. It implements a closed-loop iterative process that enables model self-correction through multi-turn interactions and tool invocation, and this pipeline can be optimized in an end-to-end manner by the GRPO algorithm.
  • Extensive evaluation of the RRVF framework across two structurally diverse domains (i.e., data charts and web interfaces) demonstrates that the proposed approach successfully learns underlying generative logic directly from pixel-level representations. The model trained via RRVF significantly outperforms state-of-the-art open-source MLLMs of comparable scale and surpasses the large-scale model used during training to provide visual feedback and judgment. Notably, it outperforms supervised fine-tuning methods on image-code pairs and exhibits superior generalization capabilities on unseen datasets. This establishes a promising paradigm for developing more capable MLLMs.
@@
Multimodal Large Language Models (MLLMs) are rapidly evolving (Liu et al. 2023; Bai et al. 2025; Zhu et al. 2025; Anthropic 2025; OpenAI 2025a; Google 2025), demonstrating remarkable capabilities in visual understanding and cross-modal interaction. At the forefront are powerful proprietary models such as GPT-o3 (OpenAI 2025a) and Gemini-2.5 (Google 2025), which employ a unified Transformer architecture for end-to-end processing of diverse inputs including images, text, and long videos. Concurrently, the open-source community is making significant strides (Team et al. 2025c; Guo et al. 2025; Team et al. 2025a). For instance, InternVL3 (Zhu et al. 2025) has achieved state-of-the-art (SOTA) performance on numerous multimodal benchmarks through its robust visual encoder and refined alignment mechanism. Similarly, models like Qwen2.5-VL (Bai et al. 2025) have shown strong capabilities in multi-language support and specialized tasks like optical character recognition (OCR) and document understanding. These collective efforts, from both proprietary and open-source domains, have significantly advanced the field and promoted the democratization of related technologies (Chu et al. 2025; Liu et al. 2025a). Despite these advances in visual representation and long-context processing, current MLLMs still have limitations in deep visual reasoning and do not fully utilize visual information (Su et al. 2025c; Wang et al. 2025c; mao2025investigatingredundancymultimodallarge).
@@
Visual reasoning, as a core capability in multimodal understanding, has garnered significant attention from the academic community in recent years (Su et al. 2025c). Efforts to bridge the semantic gap between continuous visual information and discrete language have evolved through several stages: from encoding images into static visual features for Large Language Models (LLMs), to leveraging text-based Chain-of-Thought (CoT) for reasoning (Zhang et al. 2023; Wang et al. 2025c), which explicitly breaks down complex visual problems into textual intermediate steps (Shao et al. 2024a), and more recently, to accomplishing tasks by invoking external tools (Zheng et al. 2025; Xu et al. 2025; Wu et al. 2024b; Fu et al. 2025). This paradigm, where models learn to call APIs, was pioneered in language-only contexts by works like Toolformer (Schick et al. 2023) and extended to multimodal scenarios (Wu et al. 2023). Building upon this, some models will act as advanced planners (Qi et al. 2024; Su et al. 2025b), enriching visual understanding at a finer-grained level by dynamically and iteratively invoking tools like ”zoom in” and ”crop” as reasoning steps. To further enhance the model’s autonomous cognitive abilities, subsequent research has transcended the limitations of predefined toolsets by generating executable visual code (Beltramelli 2018; Zhao et al. 2025). This approach increases reasoning flexibility by decomposing complex visual tasks into modular and programmatic sub-tasks (Wang et al. 2025a; Mallis et al. 2024; Hu et al. 2024; Shen et al. 2025b; Sur´ıs, Menon, and Vondrick 2023). Despite these advancements, robust and general-purpose multimodal reasoning remains an open challenge, with substantial room for future exploration.
@@
This section details the Reasoning-Rendering-Visual-Feedback (RRVF) framework first, then the reinforcement learning strategy for optimization.
@@
1:MLLM $\mathcal{M}$, Single Image $I$, Max Turns $T_{max}$
@@
5:  $R_{model}\leftarrow\mathcal{M}.\text{Generate}(H)$triangleright Generate response
@@
[…]
@@
[…]
12:   $F\leftarrow\text{CallTool}(C_{tool})$triangleright Invoke the external tool
13:   $H\leftarrow H\oplus F$triangleright Append visual feedback to history
@@
Table 1: Performance comparison on the ChartMimic benchmark. We report compilation success (“Exec rate”), similarity scores for specific attributes, and a quality score from “GPT-4o”, which are the metrics from the original ChartMimic benchmark (Yang et al. 2024). The best and second-best results among open-source models under 10B parameters are bolded and underlined, respectively. Results marked with * are reported by the original benchmark.
@@
Table 2: Performance comparison on the Plot2Code benchmark. The best and second-best results on the primary metrics (Exec Rate, Text, GPT-4o Score) among open-source models under 10B parameters are bolded and underlined, respectively. Results marked with * are reported by the original benchmark.
@@
$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\!\Bigl{(}\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)}\,A_{i},\,$
[…]
$-\,\beta\,D_{\mathrm{KL}}\bigl{(}\pi_{\theta}\,\big{\|}\,\pi_{\mathrm{ref}}\bigr{)}\Biggr{]},$
@@
  • Visual Similarity Reward ($R_{\text{vision}}$): This is the primary reward component, sourced directly from the Visual Judge module. It quantifies the fidelity between the final rendered image and the original input, providing a dense and accurate signal of task success. This reward can be efficiently computed using an MLLM as a judge, thereby avoiding heavy reliance on curated image-text supervision.
  • Format Correctness Reward ($R_{\text{format}}$): This component provides a crucial binary signal that encourages the generation of valid and executable code. A positive reward is granted if the MLLM’s final code is syntactically correct and can be executed by the rendering engine without errors. Otherwise, a penalty is applied. This reward discourages the model from generating nonsensical or incomplete programs.
@@
The framework’s performance is evaluated on two image-to-code tasks: chart-to-code and web-to-code. For the chart-to-code task, two standard benchmarks are utilized. The official “chart-to-code” subset of the ChartMimic dataset (Yang et al. 2024) comprises 2,400 test images. From this set, 1,800 images are repurposed for training, while the remaining 600 images, which form the official smaller test set, are used for evaluation. Additionally, the Plot2Code dataset (Wu et al. 2024a), comprising 368 test cases, is used for zero-shot evaluation. For the web-to-code task, only 1,000 screenshots are randomly sampled from the WebSight dataset (Laurençon, Tronchon, and Sanh 2024) for training, with a disjoint set of 500 images reserved for testing.
For the chart-to-code task, evaluation adheres to the official protocols and metrics of its respective benchmarks. For the web-to-code task, however, traditional metrics such as block-level HTML comparisons are ill-suited for our framework. This is because the model is designed to generate semantically equivalent but structurally distinct code, without access to the ground-truth source. Consequently, evaluation for this task is based on visual fidelity and semantic correctness, quantified using the following metrics. (1) CLIP Similarity: The perceptual similarity between the rendered image and the original input is measured by the cosine similarity of their respective CLIP embeddings (Radford et al. 2021). (2) LLM-based Assessment: A powerful multi-modal model, GPT-4o (OpenAI 2024a), is employed as an impartial judge to assess the quality of the generated outputs. It evaluates semantic consistency, element completeness, and overall visual fidelity against the input image.
@@
A comprehensive set of both closed-source and open-source models is selected for comparison to ensure a thorough evaluation. The closed-source baselines include leading proprietary models: GPT-4o (OpenAI 2024a), OpenAI O3 (OpenAI 2025a), Gemini-1.0-Pro-Vision (Google 2024), Gemini-2.5-Pro (Google 2025), and Claude-4-Sonnet (Anthropic 2025). The open-source counterparts include vision-language models such as Qwen2.5-VL (Bai et al. 2025), LLaVA-OneVision(Li et al. 2024a), Mini-Gemini (Li et al. 2024b), DeepSeek-VL (Lu et al. 2024), and InternVL3 (Zhu et al. 2025).
@@
The policy model is initialized from Qwen2.5-VL-Instruct-7B (Bai et al. 2025) and trained on a cluster of 8 NVIDIA A100 GPUs. The model is optimized using the GRPO algorithm with a global batch size of 32 and a learning rate of $1\times 10^{-6}$. During each optimization step, a group of $G=8$ candidate trajectories is sampled for each prompt. The RRVF interaction loop is set to a maximum of 4 turns, and the maximum sequence length for generation is 16,384 tokens. The KL divergence coefficient is set to 0.0.
Visual feedback and judgment are provided by a more capable model, Qwen2.5-VL-Instruct-72B (Bai et al. 2025), which serves as the reward oracle. To handle the substantial computational demand during the policy rollout phase, this judge model is deployed on a separate, dedicated cluster of 8 NVIDIA A100 GPUs. Inference is accelerated using the vLLM framework (Kwon et al. 2023) to ensure high throughput for the large volume of verification requests generated in this process.
The optimization is guided by the hybrid reward function from Equation 2. The weights are set to $w_{v}=-0.2$, $w_{f}=0.8$, and $w_{t}=1.0$. The interaction loop operates for a maximum of $T_{max}=4$ turns. Within this loop, the tool-use efficiency reward, $R_{\text{eff}}$, assigns a reward of 1/3 for each successful tool call and is fixed to its maximum value of 1.0 if the visual similarity score surpasses 0.95.
@@
During the evaluation phase, the model operates in a direct, single-turn inference mode. The final output is generated from the input prompt in a single pass, without access to the iterative refinement mechanism or any external tools available during the training loop. This protocol is intentionally designed to rigorously assess the model’s internalized multi-modal comprehension and generation capabilities, demonstrating the improvements gained autonomously after the completion of training.
@@
The main experimental results are presented in Table 1, Table 2, and Table 3. The findings indicate that the RRVF-trained model achieves state-of-the-art performance among open-source Multimodal Large Language Models (MLLMs) of comparable size. This consistent outperformance across diverse tasks highlights the efficacy and generalizability of the proposed framework, which learns complex visual-to-code logic solely from visual feedback.
@@
On the Plot2Code dataset (Table 2), the RRVF model achieves the highest execution rate (96.21%) among all open-source competitors. This outstanding reliability is directly reflected in the primary metrics (Text and GPT-4o Score), which penalize execution failures by being multiplied by the execution rate. On these realistic metrics, RRVF surpasses all similar-sized models, underscoring its enhanced robustness and reliability. When considering “pass” scores, which are calculated only on successfully executed code, models with lower execution rates are effectively evaluated on easier subsets of data. Despite RRVF’s high execution rate exposing it to more challenging examples, its performance remains highly competitive.
@@
The applicability of RRVF is further tested on the WebSight benchmark, a domain characterized by high structural freedom (Table 3). In this task, our RRVF-trained model achieves strong performance with a CLIP score of 88.29 and a GPT score of 91.50, surpassing the similar-sized models. Crucially, this advantage is obtained without any access to ground-truth HTML, relying instead on our proposed method. This demonstrates that the framework can effectively learn to deconstruct and replicate complex layouts and component relationships from raw pixels alone, validating the potential of learning through pure visual feedback even in highly unstructured domains.
@@
Our ablation study, conducted on the ChartMimic dataset, directly pits RRVF against a standard Supervised Fine-Tuning (SFT) baseline. Although SFT benefits from direct access to ground-truth code, RRVF demonstrates marked superiority. As shown in Table 1, RRVF achieves a near-perfect execution rate of 97.83%, massively outperforming SFT’s 69.00%. This highlights the reliability of the code generated by our framework. Furthermore, RRVF’s dominant Layout score (80.97 vs. 63.60) confirms its advanced comprehension of holistic visual structures. This indicates that learning from the final visual outcome, as RRVF does, cultivates a more robust and compositional reasoning ability compared to the surface-level mimicry of SFT. A more profound finding is that our final 7B model, trained via RRVF, surpasses its much larger 72B teacher model in the Overall score (64.36 vs. 47.30). This counter-intuitive result proves that RRVF is more than knowledge distillation; it enables the policy model to actively explore the solution space and discover policies that are superior to its own feedback provider, thus bootstrapping its capabilities beyond initial constraints.
@@
To evaluate the true extent of the learned skills, we test the models’ generalization ability via a zero-shot evaluation on the unseen Plot2Code dataset. The SFT-trained model exhibits a classic case of overfitting; its execution rate dramatically plummets from 69.00% on the training domain to a mere 49.24% on the new dataset. This sharp decline suggests that SFT learns brittle, non-transferable knowledge tied to the specific code patterns of the ChartMimic dataset. In stark contrast, our RRVF-trained model showcases exceptional generalization. Its execution rate remains remarkably stable, decreasing only negligibly from 97.83% to 96.21%. This stability is strong evidence that RRVF learns the fundamental and transferable principles of translating visual elements into programmatic logic, rather than simply memorizing code templates. This result underscores a key advantage of our framework: it fosters robust reasoning skills that reliably generalize to new and unfamiliar challenges.

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) exhibit impressive performance across various visual tasks. Subsequent investigations into enhancing their visual reasoning abilities have significantly expanded their performance envelope. However, a critical bottleneck in the advancement of MLLMs toward deep visual reasoning is their heavy reliance on curated image-text supervision. To solve this problem, we introduce a novel framework, “Reasoning-Rendering-Visual-Feedback” (RRVF), that enables MLLMs to learn complex visual reasoning from only raw images. This framework builds on the “Asymmetry of Verification” principle, i.e., verifying the rendered output against the source image is substantially easier than performing deep visual reasoning to generate a faithful, structured representation such as code. We demonstrate that this relative ease provides an ideal reward signal for optimization via Reinforcement Learning (RL), thereby reducing reliance on image-text supervision. RRVF implements a closed-loop iterative process encompassing reasoning, rendering, and visual feedback components, enabling the model to perform complex reasoning, including self-correction through multi-turn interactions. This process is optimized end-to-end using the GRPO algorithm. Extensive evaluations are conducted on image-to-code generation across two diverse domains: data charts and web interfaces. The RRVF-trained model not only outperforms existing similarly sized open-source MLLMs and supervised fine-tuning baselines but also exhibits superior generalization. Notably, the model outperforms the more advanced MLLM used to generate visual feedback during training. Code is available at https://github.com/L-O-I/RRVF.
@@ Introduction
Figure 1: Comparison of training paradigms. Conventional MLLMs (top) rely on paired image-text data. Our RRVF (bottom) trains solely on raw images via a closed-loop reasoning-rendering-visual-feedback process under GRPO.
Figure 2: RRVF is a training framework that improves visual reasoning ability using only image inputs. Given an input image (a), the multimodal language model performs iterative reasoning (b) by generating rendering code. The code is executed by external tools, and the output is compared with the original image by a visual judge (d). The comparison results are converted into structured feedback (c), which guides the next round of code generation. The reward manager (e) combines visual similarity, format validity, and tool usage efficiency. These signals are used to optimize the model through reinforcement learning, implemented with the GRPO algorithm.
@@
Specifically, for MLLMs, effective reasoning necessitates capabilities that extend beyond elementary visual perception to achieve sophisticated visual comprehension and inference. This paradigm requires models to not only perform object recognition but also to analyze underlying visual semantics, spatial-geometric relationships, and complex inter-object dependencies (Su et al. 2025c; Luo et al. 2024). While accurate visual perception serves as a fundamental prerequisite for such reasoning capabilities, it inherently depends on robust reasoning mechanisms, creating a challenging bidirectional dependency where perception and reasoning are mutually reinforcing.
@@
Moreover, the GRPO optimization is guided by a hybrid reward function. (1) The primary component is a visual similarity reward that quantifies the fidelity between the rendered image and the original image by integrating signals from established vision-language models. (2) This is supplemented by a format correctness reward to ensure the generation of syntactically correct and executable code. (3) An adaptive tool-use reward is introduced to balance exploration and convergence. This hybrid reward function effectively enables the model to learn underlying generative logic directly from pixel-level visual representations.
@@
  • The exploration of the “Asymmetry of Verification” principle for MLLM training in complex visual reasoning tasks. We establish that verifying the visual correctness of a rendered output against a source image is a significantly more tractable problem than generating the initial code. This asymmetry provides an effective learning signal that circumvents the need for textual supervision.
  • A novel framework, termed Reasoning-Rendering-Visual-Feedback (RRVF), is proposed. It implements a closed-loop iterative process that enables model self-correction through multi-turn interactions and tool invocation, and this pipeline is optimized in an end-to-end manner using the GRPO algorithm.
  • Extensive evaluation of the RRVF framework across two structurally diverse domains (i.e., data charts and web interfaces) demonstrates that the proposed approach successfully learns underlying generative logic directly from pixel-level representations. The model trained with RRVF significantly outperforms state-of-the-art open-source MLLMs of comparable scale as well as supervised fine-tuning approaches on image-code pairs, exhibiting superior generalization on unseen datasets. Notably, it also surpasses the larger model used during training to provide visual feedback and judgment. This establishes a promising paradigm for developing more capable and robust MLLMs.
@@
Multimodal Large Language Models (MLLMs) are rapidly evolving (Liu et al. 2023; Bai et al. 2025; Zhu et al. 2025; Anthropic 2025; OpenAI 2025a; Google 2025), demonstrating remarkable capabilities in visual understanding and cross-modal interaction. At the forefront are powerful proprietary models such as OpenAI o3 (OpenAI 2025a) and Gemini-2.5 (Google 2025), which employ a unified Transformer architecture for end-to-end processing of diverse inputs, including images, text, and long videos. Concurrently, the open-source community is making significant strides (Team et al. 2025c; Guo et al. 2025; Team et al. 2025a). For instance, InternVL3 (Zhu et al. 2025) has achieved state-of-the-art (SOTA) performance on numerous multimodal benchmarks through its robust visual encoder and refined alignment mechanism. Similarly, models like Qwen2.5-VL (Bai et al. 2025) have shown strong capabilities in multi-language support and specialized tasks like optical character recognition (OCR) and document understanding. These collective efforts, from both proprietary and open-source domains, have significantly advanced the field and promoted the democratization of related technologies (Chu et al. 2025; Liu et al. 2025a). Despite these advances in visual representation and long-context processing, current MLLMs still have limitations in deep visual reasoning and do not fully utilize visual information (Su et al. 2025c; Wang et al. 2025c; Mao et al. 2025).
@@
Visual reasoning, as a core capability in multimodal understanding, has garnered significant attention from the academic community in recent years (Su et al. 2025c). Efforts to bridge the semantic gap between continuous visual information and discrete language have evolved through several stages: from encoding images into static visual features for Large Language Models (LLMs), to leveraging text-based Chain-of-Thought (CoT) for reasoning (Zhang et al. 2023; Wang et al. 2025c), which explicitly breaks down complex visual problems into textual intermediate steps (Shao et al. 2024a), and more recently, to accomplishing tasks by invoking external tools (Zheng et al. 2025; Xu et al. 2025; Wu et al. 2024b; Fu et al. 2025). This paradigm, where models learn to call APIs, was pioneered in language-only contexts by works like Toolformer (Schick et al. 2023) and extended to multimodal scenarios (Wu et al. 2023). Building upon this, some models act as advanced planners (Qi et al. 2024; Su et al. 2025b), enriching visual understanding at a finer-grained level by iteratively invoking tools like “zoom in” and “crop” as reasoning steps. To further enhance the model’s autonomous cognitive abilities, subsequent research has transcended the limitations of predefined toolsets by generating executable visual code (Beltramelli 2018; Zhao et al. 2025; Wang et al. 2025a; Shen et al. 2025b; Wu et al. 2025; Xiao et al. 2025; Liu et al. 2025a). Despite these advancements, achieving robust and powerful multimodal reasoning remains an open challenge.
@@
This section first details the Reasoning-Rendering-Visual-Feedback (RRVF) framework and then describes the reinforcement learning strategy used for its optimization.
@@
1:MLLM $\mathcal{M}$, Single Image II, Max Turns $T_{max}$
@@
5:  $R_{model}\leftarrow\mathcal{M}.\text{Generate}(H)$ $\triangleright$ Generate response
@@
8:   $C_{final}\leftarrow\text{Extract}(R_{model},\texttt{<answer>})$
@@
11:   $C_{tool}\leftarrow\text{Extract}(R_{model},\texttt{<tool\_call>})$
12:   $F\leftarrow\text{CallTool}(C_{tool})$ $\triangleright$ Invoke the external tool
13:   $H\leftarrow H\oplus F$ $\triangleright$ Append visual feedback to history
@@
Table 1: Performance comparison on the ChartMimic benchmark. We report the metrics from the original ChartMimic benchmark (Yang et al. 2024). The best and second-best results among open-source models under 10B parameters are bolded and underlined, respectively. Results marked with * are reported by the original benchmark.
@@
Table 2: Performance comparison on the python_matplotlib subset of the Plot2Code benchmark. The best and second-best results on the primary metrics (Exec Rate, Text, GPT-4o Score) among open-source models under 10B parameters are bolded and underlined, respectively. Results marked with * are reported by the original benchmark.
@@
$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\Bigl{(}\min\!\Bigl{(}\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)}\,A_{i},\,$
$\mathrm{clip}\!\Bigl{(}\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)},1-\varepsilon,1+\varepsilon\Bigr{)}A_{i}\Bigr{)}$
$-\,\beta\,D_{\mathrm{KL}}\bigl{(}\pi_{\theta}\,\big{\|}\,\pi_{\mathrm{ref}}\bigr{)}\Bigr{)}\Biggr{]},$
@@
  • Visual Similarity Reward ($R_{\text{vision}}$): This is the primary reward component, sourced directly from the Visual Judge. It quantifies the fidelity between the final rendered image and the original input, providing a dense and accurate signal. This reward can be efficiently computed using an MLLM as a judge, thereby avoiding heavy reliance on curated image-text supervision.
  • Format Correctness Reward ($R_{\text{format}}$): This reward penalizes structural and syntactic errors. A penalty is applied if the reasoning, tool_call, and answer steps are improperly formatted or if the generated code is non-executable.
@@
The framework’s performance is evaluated on two image-to-code tasks: chart-to-code and web-to-code. For the chart-to-code task, two standard benchmarks are utilized. The official “chart-to-code” subset of the ChartMimic dataset (Yang et al. 2024) comprises 2,400 test images. From this set, 1,800 images are repurposed for training, while the remaining 600 images, which form the official smaller test set, are used for evaluation. Additionally, zero-shot evaluation is performed on the python_matplotlib subset of the Plot2Code dataset (Wu et al. 2024a), which comprises 132 test cases. For the web-to-code task, only 2,000 screenshots are randomly sampled from the WebSight dataset (Laurençon, Tronchon, and Sanh 2024) for training, with a disjoint set of 500 images reserved for testing.
For the chart-to-code task, evaluation adheres to the official protocols and metrics of their respective benchmarks. For the web-to-code task, however, traditional metrics such as block-level HTML comparisons are ill-suited for our framework. This is because the model is designed to generate semantically equivalent but structurally distinct code, without access to the ground-truth source. Consequently, evaluation for this task is based on visual fidelity and semantic correctness, quantified using the following metrics. (1) CLIP Similarity: The perceptual similarity between the rendered image and the original input is measured by the cosine similarity of their respective CLIP embeddings (Radford et al. 2021). (2) LLM-based Assessment: A powerful multimodal model, GPT-4o (OpenAI 2024a), is employed as an impartial judge to assess the quality of the generated outputs. It evaluates semantic consistency, element completeness, and overall visual fidelity against the input image.
@@
A comprehensive set of both closed-source and open-source models is selected for comparison to ensure a thorough evaluation. The closed-source baselines include leading proprietary models: GPT-4o (OpenAI 2024a), OpenAI o3 (OpenAI 2025a), Gemini-2.5-Pro (Google 2025), and Claude-4-Sonnet (Anthropic 2025). The open-source counterparts include vision-language models such as LLaVA-OneVision (Li et al. 2024), Qwen2.5-VL (Bai et al. 2025), and InternVL3 (Zhu et al. 2025).
@@
The policy model is initialized from Qwen2.5-VL-Instruct-7B (Bai et al. 2025) and trained on a cluster of 8 NVIDIA A100 GPUs. The model is optimized using the GRPO algorithm with a global batch size of 32 and a learning rate of $1\times 10^{-6}$. During each optimization step, a group of $G=8$ candidate trajectories is sampled for each prompt. The RRVF interaction loop is set to a maximum of 4 turns, and the maximum sequence length for generation is 16,384 tokens. The KL divergence coefficient is set to $0.$
Visual feedback and judgment are provided by a more capable model, Qwen2.5-VL-Instruct-72B (Bai et al. 2025), which serves as the reward oracle. To handle the substantial computational demand during the policy rollout phase, this judge model is deployed on a separate, dedicated cluster of 8 NVIDIA A100 GPUs. Inference is accelerated using the vLLM framework (Kwon et al. 2023).
The optimization is guided by the hybrid reward function from Equation 2. The weights are set to $w_{v}=0.2$, $w_{f}=0.8$, and $w_{t}=1.0$. The interaction loop operates for a maximum of $T_{max}=4$ turns. Within this loop, the tool-use efficiency reward, $R_{\text{tool}}$, assigns a reward of 1/3 for each successful tool call and is fixed to its maximum value of 1.0 if the visual similarity score surpasses 0.95.
Details of the cold-start and the utilized prompts are provided in Appendix A and Appendix C, respectively.
@@
During the evaluation phase, the model operates in a direct, single-turn inference mode. The final output is generated from the input prompt in a single pass, without access to the iterative refinement mechanism or any external tools available during the training loop. This protocol is intentionally designed to rigorously assess the model’s internalized multimodal comprehension and generation capabilities, demonstrating the improvements gained autonomously after the completion of training.
@@
The main experimental results are presented in Table 1, Table 2, and Table 3. The findings indicate that the RRVF-trained model achieves SOTA performance among open-source MLLMs of comparable size. Case studies are provided in Appendix B.
@@
On the Plot2Code dataset (Table 2), the RRVF model achieves the highest execution rate (96.21%) among all models. Consequently, RRVF surpasses similarly sized models on the primary Text and GPT-4o Score metrics, underscoring its enhanced robustness and reliability. When considering “pass” scores, which are calculated only on successfully executed code, models with lower execution rates are effectively evaluated on easier subsets of data. Despite RRVF’s high execution rate exposing it to more challenging examples, its performance remains highly competitive.
@@
The applicability of RRVF is further tested on the WebSight benchmark, a domain characterized by high structural freedom (Table 3). In this task, our RRVF-trained model achieves strong performance with a CLIP score of 88.29 and a GPT score of 91.50, surpassing similarly sized models. Crucially, this advantage is obtained without any access to ground-truth HTML, relying instead on our proposed method. This demonstrates that the framework can effectively learn to deconstruct and replicate complex layouts and component relationships from raw pixels alone, validating the potential of learning through pure visual feedback even in highly unstructured domains.
Training Analysis
Figure 3: Training curves for reward score and tool usage.
Figure 3 illustrates the training process, where a steadily increasing reward score confirms policy improvement. The number of tool calls follows a distinct dip-rise-plateau trend, aligning with trends reported in prior work (Zheng et al. 2025), reflecting the agent’s progression from learning the tool’s mechanics to fully exploiting iterative feedback.
@@
Our ablation study, conducted on the ChartMimic dataset, directly pits RRVF against a standard Supervised Fine-Tuning (SFT) baseline. Although SFT benefits from direct access to ground-truth code, RRVF demonstrates marked superiority. As shown in Table 1, RRVF achieves a near-perfect execution rate of 97.83%, significantly outperforming SFT’s 69.00%. This highlights the reliability of the code generated by our framework. Furthermore, RRVF’s dominant Layout score (80.97 vs. 63.60) confirms its advanced comprehension of holistic visual structures. This indicates that learning from the final visual outcome, as RRVF does, cultivates a more robust and compositional reasoning ability compared to the surface-level mimicry of SFT. A more profound finding is that our final 7B model, trained via RRVF, surpasses its much larger 72B teacher model in the Overall score (64.36 vs. 47.30). This counter-intuitive result proves that RRVF is more than knowledge distillation; it enables the policy model to actively explore the solution space and discover policies that are superior to its own feedback provider, thus bootstrapping its capabilities beyond initial constraints.
@@
To evaluate the true extent of the learned skills, we test the models’ generalization ability via a zero-shot evaluation on the unseen Plot2Code dataset. As shown in Table 2, the SFT-trained model’s execution rate dramatically plummets from 69.00% on the Chartmimic to 49.24% on the Plot2Code. This sharp decline suggests that SFT learns brittle, non-transferable knowledge tied to the specific code patterns of the ChartMimic dataset. In stark contrast, our RRVF-trained model showcases exceptional generalization. Its execution rate remains remarkably stable, decreasing only negligibly from 97.83% to 96.21%. This stability is strong evidence that RRVF learns the fundamental and transferable principles of translating visual elements into programmatic logic, rather than simply memorizing code templates.